OpenStudy (anonymous):

Find the dimensions of the rectangle where the length is 5 meters less than twice its width and whose area is 35 m^2

terenzreignz (terenzreignz):

This'll be slightly trickier. First, let w be your width, what's your length?

OpenStudy (anonymous):

length: 5-2w? right?

terenzreignz (terenzreignz):

No... Twice its width is 2w Five meters less twice its width is 5 - 2w Five meters less THAN twice its width is 2w - 5

OpenStudy (anonymous):

Oh, okay so Width:w Length: 2w-5 Area: 35^2

terenzreignz (terenzreignz):

Area is also length times width. Length times width is w(2w - 5) = 2w² - 5w

terenzreignz (terenzreignz):

And this is equal to 35. Care to solve for w? :) 2w² - 5w = 35

OpenStudy (chihiroasleaf):

solve the quadratic equation... have you learned about solving quadratic equation? :)

OpenStudy (anonymous):

Yes, i just always have a hard time solving them. Im so terrible at most thing in math:(

OpenStudy (chihiroasleaf):

well.., we have to solve the quadratic equation \[ \Large 2w^2 -5w = 35\] it's easier to make one side become zero.., so we have \[ \Large 2w^2 -5w - 35 = 0\] do you quadratic formula?